OpenStudy (anonymous):

The length of a rectangle is 12 m, and its width is 16 m. What is the area of half of the rectangle?

OpenStudy (amistre64):

which half?

OpenStudy (anonymous):

its 96 right

OpenStudy (anonymous):

same as the the area of the triangle with base 12 and height 16

OpenStudy (amistre64):

12*16 ----- = 12*8 = 96 yes 2

OpenStudy (anonymous):

yes that's right :)

OpenStudy (anonymous):

thanks XD

OpenStudy (anonymous):

Well done sunset!

OpenStudy (anonymous):

6* 16 = 96